# SignRoundV2: Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs

### v1

@@ 3.4 Hyperparameters and Tricks
16 16bits 64.16 70.00 65.67 63.24 67.00
@@
Table 3: Comparison with head-layer (Head)/ tail-layer (Tail) heuristics at 4.5-bit and 5-bit MXFP with only DL, without tuning. The average accuracies are across 10 tasks (detailed in Section 4.3). The "I" in the model name is short for "Instruct".
Avg. bits Method Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
4.5 Head 8-bit 59.18 60.63 58.92
Tail 8-bit 58.93 61.38 59.26
DL 60.64 63.13 59.81
5 Head 8-bit 60.82 61.75 59.43
Tail 8-bit 60.12 61.35 59.65
DL 62.57 64.04 61.45
6 Head 8-bit 59.88 62.22 60.54
Tail 8-bit 62.65 63.39 61.53
DL 63.64 65.45 62.00
Table 4: The average accuracies across 10 tasks(Section 4.3) at an average of 3 bits (W2G128 / W4G128). The "I" in the model name is short for "Instruct".
Avg. bits = 3 Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
Head 4-bit 31.98 32.70 31.96
Tail 4-bit 60.58 37.98 45.36
Ours 61.48 40.58 48.62
4 Experiments
4.1 Experimental Setup
Models and Benchmarks.
We evaluate SignRoundV2 on two major LLM families, LLaMA (touvron2023llamav2; grattafiori2024llama) and Qwen (bai2023qwen; yang2025qwen3), under both low-bit and mixed-precision quantization settings across standard benchmarks. All evaluations follow the lm-eval-harness framework (eval-harness). The benchmark suite includes ARC-Challenge (clark2018think), ARC-Easy (clark2018think), BoolQ (clark2019boolq), HellaSwag (Zellers_2019), LAMBADA (paperno2016lambada), MMLU (hendrycks2020measuring), OpenBookQA (mihaylov2018can), PIQA (bisk2020piqa), TruthfulQA (lin2022truthfulqa), and WinoGrande (sakaguchi2021winogrande).
Quantization Configurations.
Following SignRoundV1 (cheng2024optimize), we use The Pile (gao2020pile) as the calibration dataset and adopt symmetric quantization. All experiments are conducted on NVIDIA A100 80GB GPUs. For simplicity, we omit the bit costs associated with scales and zero points. For MXFP4 and MXFP8, we follow the standard definitions in (mxfp4).
4.2 Comparison of Results at W2A16/W4A16
Table 1 presents the average accuracies across 5 tasks (ARC-Challenge, ARC-Easy, HellaSwag, PIQA and WinoGrande) for W2A16 with and without mixed-precision quantization on various LLaMA models. At pure W2A16 setting, our method consistently outperforms existing PTQ approaches such as GPTQ  (frantar2022gptq), AWQ  (lin2024awq), OmniQuant  (shao2023omniquant), and SignRoundV1  (cheng2024optimize) by a substantial margin. Compared to high cost methods like AQLM  (egiazarian2024extreme), QuIP# (tseng2024quip), and EfficientQAT  (chen2025efficientqat), our approach achieves comparable performance on large models, while slightly trailing on smaller models. At slightly higher bit widths (e.g., 2.5 bits), with the support of DeltaLoss, our method continues to deliver strong accuracy, demonstrating the effectiveness of adaptive layer-wise quantization. Overall, these results confirm that our approach reliably preserves model quality under extreme low-bit mixed-precision settings. Detailed accuracy results for each task are provided in Appendix A.
4.3 Comparison of Results at MXFP4/8
Table 2 reports the average accuracies and recovery rates across 10 tasks (all tasks listed in Section 4.1) for MXFP4/8 mixed-precision quantization. Since few prior works report results for standard MXFP4, we primarily compare against our own baselines. The results show that our method consistently outperforms RTN and SignRoundV1 (SRV1) at low average bit widths (4-5 bits), achieving higher accuracy and recovery rates across nearly all evaluated models. At 4 bits, our approach improves recovery by $\sim$1-3% compared to SRV1. At slightly higher bit widths (5-6 bits), both DeltaLoss-only (DL) and our method approach full-precision accuracy, but our method achieves the highest recovery in most scenarios, reaching near or above 99% across all models. Overall, these results demonstrate that our approach effectively preserves model performance under extreme low-bit mixed-precision quantization while minimizing accuracy loss. Detailed accuracy results for each task are provided in Appendix A.
4.4 Comparison of Mixed Bits Recipes
The results in Tables 3 and 4 demonstrate that our method consistently outperforms simple head-layer and tail-layer heuristics across a range of precision budgets and model architectures. For MXFP quantization (4.5, 5, and 6 bits), the DL strategy achieves the highest average accuracy for all three models without requiring any tuning, showing clear improvements over allocating 8-bit precision only to the head (near the LM head) or tail layers (near the embedding). At lower precision (average 3 bits using W2G128/W4G128), the performance gap becomes even more obvious: both head- and tail-focused heuristics degrade substantially, while our method remains robust and delivers the best accuracy across all evaluated models.
4.5 Ablation Study of Initialization
We evaluate the impact of our initialization strategy on Qwen3-8B and Llama3.1-8B-Instruct under the W2A16G64 setting on SignRoundV1 with or without initialization. As shown in Table 5, enabling initialization consistently improves accuracy across all five tasks. Overall, the results demonstrate that proper pre-tuning initialization significantly stabilizes optimization and leads to substantial accuracy improvements for both models.
Table 5: Ablation study of pre-tuning initialization for Qwen3-8B and Llama3.1-8B-Instruct at W2A16G64.
Model Method Arc-c hella. lamb. mmlu wino.
Qwen3-8B W/O init 34.90 44.87 47.18 54.09 63.85
With init 43.69 45.12 54.61 56.12 66.22
Llama3.1-8B-I W/O init 36.43 46.20 52.57 48.05 65.67
With init 36.43 47.32 60.41 50.11 67.40
4.6 Quantization Cost
Table 6 compares the end-to-end quantization time across several representative methods. Traditional QAT-based and vector quantization approaches such as EfficientQAT (chen2025efficientqat), QuIP#  (tseng2024quip), and AQLM  (egiazarian2024extreme) require tens to hundreds of hours to converge. Besides, we argue that QAT requires careful hyperparameter tuning, and its actual quantization cost is often much higher than what is reported. In contrast, SignRoundV1 completes in only 2.2 hours. Our method maintains a similarly low computational cost, requiring 2.5 hours, while the enhanced version (Ours*) takes 6 hours due to additional tuning samples and steps. Overall, our approach achieves competitive accuracy with substantially lower time overhead compared with prior methods.
Table 6: Time Cost Comparison for Llama2-70B
Method One A100-80GB? GPU hours
SignRoundV1 2.2
EfficientQAT 41
QuIP# × 270
AQLM 336
Ours 2.5
Ours* 6
Table 7 summarizes the VRAM and runtime overhead of computing DeltaLoss for weight-only quantization (e.g., W2A16), with MXFP4 exhibiting comparable cost. The results show that DeltaLoss remains computationally practical even for 70B-scale models.
Table 7: VRAM and Time Cost of DeltaLoss For W2A16
Model VRAM Cost (GB) Time Cost (s)
Qwen3-8B 15 $60\times len\ of\ options$
Qwen3-32B 30 $180\times len\ of\ options$
Llama-2-70B 40 $420\times len\ of\ options$
5 Conclusion
We introduced SignRoundV2, a practical framework for extreme low-bit post-training quantization of LLMs. By combining a gradient-informed sensitivity metric with a lightweight pre-tuning search, SignRoundV2 enables stable and accurate quantization even at aggressive bit widths. Experiments show that it delivers production-grade accuracy at 4–5 bits and strong performance at 2 bits. Our results demonstrate that effective PTQ can significantly reduce the deployment cost of LLMs while preserving model quality.
6 Limitations
Despite its strong performance, SignRoundV2 has several limitations. First, without mixed precision, it still exhibits a noticeable gap from the full-precision model at extremely low bit widths, particularly for smaller models. Second, the bit configuration is fixed before tuning, without considering how much quantization error could be mitigated during the tuning process. Third, the method requires gradient computation, which limits its applicability in frameworks such as ONNX that do not support gradient-based optimization.
Ethics Statement
Our research aims to advance knowledge in LLM quantization. SignRoundV2 utilizes open-source models and publicly available datasets, and is not tied to particular applications, requiring only minimal fine-tuning steps on the original models. This ensures that the technical details of our method carry no potential ethical implications. We acknowledge the contributions of the creators and maintainers of these resources and provide citations to the original sources.
Appendix A Detailed Data
This appendix provides detailed accuracy results for each task.

### v2

@@ 4.4 Performance under MXFP Quantization
16 16-bit 64.16 70.00 65.67 63.24 67.00
4 RTN 58.31 (90.88%) 68.76 (98.23%) 60.62 (92.32%) 58.54 (92.57%) 64.87 (96.82%)
Ours 61.47 (95.81%) 69.31 (99.01%) 63.72 (97.03%) 61.45 (97.17%) 65.90 (98.36%)
4.5 DL 62.00 (96.63%) 70.17 (100.24%) 64.16 (97.70%) 61.35 (97.01%) 66.26 (98.90%)
Ours 63.34 (98.72%) 70.31 (100.44%) 64.77 (98.63%) 62.74 (99.21%) 67.10 (100.15%)
5 DL 63.53 (99.02%) 70.69 (100.99%) 65.22 (99.31%) 62.06 (98.13%) 66.41 (99.12%)
Ours 64.00 (99.75%) 70.66 (100.94%) 65.35 (99.51%) 62.81 (99.32%) 67.19 (100.28%)
4.5 Ablation Studies
We conduct ablation studies to isolate the individual contributions of pre-tuning initialization and loss filtering. As summarized in Table 4, both strategies consistently enhance performance under the W2A16G64 setting. Notably, incorporating loss filtering yields a significant boost in average accuracy (+0.82% on Qwen3), while its synergy with proper initialization (Ours) achieves a substantial 3.41% absolute improvement on Llama3.1. These results underscore that stabilizing the initial optimization phase is paramount for preserving the reasoning capabilities of ultra-low-bit quantized models.
Table 4: Ablation studies of loss filtering and pre-tuning initialization at W2A16G64.
Model Method ARC-c ARC-e BoolQ Hella LAMB MMLU OBQA PIQA Truth. Wino AVG.
Qwen3-8B V1 38.65 73.11 80.83 44.60 48.28 56.19 28.40 71.82 30.84 63.14 53.59
V1+Filtering 45.14 75.97 78.62 44.78 51.50 55.23 25.80 72.36 31.21 64.48 54.51
V1+Init 40.78 73.82 76.94 44.75 51.29 56.77 28.60 71.33 32.07 64.25 54.06
Ours 43.52 75.13 74.89 45.10 52.80 57.56 26.00 71.22 31.70 64.64 54.26
Llama3.1-8B-I V1 36.95 69.82 73.06 44.89 50.15 43.80 25.20 73.18 29.74 65.19 51.20
V1+Filtering 36.86 69.99 70.89 46.72 55.48 46.05 27.80 72.31 31.58 65.75 52.34
V1+Init 36.52 71.76 80.21 46.87 59.11 48.28 26.40 72.91 29.74 66.69 53.85
Ours 38.65 71.21 77.46 47.27 60.45 49.67 26.20 73.61 29.87 67.48 54.19
4.6 Quantization Cost
Table 6 compares end-to-end quantization time. QAT-based and vector quantization methods (e.g., EfficientQAT, QuIP#, AQLM) typically require tens to hundreds of GPU hours and careful hyperparameter tuning, often leading to higher practical costs than reported. In contrast, SignRoundV1 completes in 2.2 hours, while our method requires 2.5 hours, and 6 hours for the enhanced version (Ours*). Overall, our approach achieves competitive accuracy with substantially lower time cost.
Table 6 reports the additional VRAM and runtime overhead of DeltaLoss for weight-only quantization (e.g., W2A16), which remains comparable to MXFP4.
Table 5: Runtime Cost Comparison for Llama2-70B
Method Fits A100-80GB GPU hours
SignRoundV1 $\checkmark$ 2.2
EfficientQAT $\checkmark$ 41
QuIP# $\times$ 270
AQLM $\checkmark$ 336
Ours $\checkmark$ 2.5
Ours* $\checkmark$ 6
Table 6: Cost Comparison for Adaptive Bit-Width Assignment (Excluding Quantization Overhead).
Model Method VRAM (GB) GPU hours
Qwen3-8B AMQ 20 7
SFMP 48 0.05
Ours 15 $0.017{\times}\text{len(opt)}$
Qwen3-32B AMQ 70 $\approx 48$
SFMP 150 $\approx 0.2$
Ours 30 $0.05{\times}\text{len(opt)}$
Llama2-70B AMQ 160 176
SFMP 350 0.6
Ours 40 $0.117{\times}\text{len(opt)}$
5 Conclusion
We present SignRoundV2, a robust post-training quantization framework engineered for extremely low-bit LLM deployment. By integrating a gradient-informed sensitivity metric with a lightweight pre-tuning mechanism, SignRoundV2 ensures stable and accurate quantization under aggressive compression regimes. Extensive experiments demonstrate that our approach maintains high-fidelity performance at 4–5 bits and yields robust results even at 2 bits. These results confirm that effective PTQ can significantly reduce the computational and memory footprints of LLMs, offering a scalable solution for efficient large-scale model inference.
6 Limitations
Despite its strong performance, SignRoundV2 has several limitations. First, in uniform 2-bit settings, it still exhibits a noticeable accuracy gap compared with full-precision baselines, indicating that extremely low-bit PTQ remains challenging without mixed precision or additional training. Second, the bit-width configuration is determined before block-wise tuning and remains static during optimization; jointly updating bit allocation and rounding parameters may further improve accuracy but would increase search complexity. Third, DeltaLoss relies on gradient computation, which limits its immediate applicability in inference-only frameworks such as ONNX Runtime.
Ethics Statement
This research aims to advance efficient LLM quantization. SignRoundV2 leverages open-source models and publicly available datasets, adhering to their respective licenses and usage terms. As our method requires only minimal tuning of pre-existing models and is not tied to specific high-risk applications, it carries no significant new ethical risks. We acknowledge the contributions of the creators and maintainers of these resources and provide proper citations to the original sources.
Appendix A Comparison with Naive Mixed-bit Methods
The results in Tables 7 and 8 demonstrate that our method consistently outperforms simple head-layer and tail-layer heuristics across a range of precision budgets and model architectures. For MXFP quantization (4.5, 5, and 6 bits), the DL strategy achieves the highest average accuracy for all three models without requiring any tuning, showing clear improvements over allocating 4-bit precision only to the head (near the LM head) or tail layers (near the embedding). At lower precision (average 3 bits using W2G128/W4G128), the performance gap becomes even more obvious: both head- and tail-focused heuristics degrade substantially, while our method remains robust and delivers the best accuracy across all evaluated models.
Table 7: Comparison with head-layer (Head) and tail-layer (Tail) heuristics at 4.5, 5, and 6 bits using DeltaLoss without tuning. The average accuracies are computed over 10 tasks; see Section 4.4 for details. The “I” in the model name is short for “Instruct”.
Avg. bits Method Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
4.5 Head 4-bit 58.93 61.38 59.26
Middle 8-bit 58.96 60.79 59.36
Middle 4-bit 58.94 62.21 59.50
Tail 4-bit 59.18 60.63 58.92
DL 60.64 63.13 59.81
5 Head 4-bit 60.12 61.35 59.65
Middle 8-bit 59.40 61.03 60.26
Middle 4-bit 59.61 62.79 60.20
Tail 4-bit 60.82 61.75 59.43
DL 62.57 64.04 61.45
6 Head 4-bit 62.65 63.39 61.53
Middle 8-bit 60.92 62.87 61.42
Middle 4-bit 61.80 62.91 60.65
Tail 4-bit 59.88 62.22 60.54
DL 63.64 65.45 62.00
Table 8: Average accuracies across 10 tasks at an average of 3 bits (W2G128 / W4G128). The “I” in the model name is short for “Instruct”.
Avg. bits = 3 Llama3.1-8B-I Qwen2.5-7B-I Qwen3-8B
Head 4-bit 31.98 32.70 31.96
Tail 4-bit 60.58 37.98 45.36
Ours 61.48 40.58 48.62
Appendix B Comparison with Prior Mixed-bit Methods
This appendix provides the detailed six-task accuracies for the mixed-bit weight-only quantization results reported in Section 4.3. We report HellaSwag, WinoGrande, ARC-Easy, ARC-Challenge, PIQA, BoolQ, and the average accuracy (AVG) for each model and method. Gray rows denote uniform-bit baselines, namely W3G128 and W4G128. Specifically, Table 9 reports results for Llama-3.1-8B, Table 10 for Qwen3-8B, Table 11 for Qwen3-14B, and Table 12 for Qwen3-32B.
Table 9: Detailed accuracies on six tasks for Llama3.1-8B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Gray rows denote uniform-bit baselines.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 54.78 82.70 83.12 79.26 80.90 74.19 75.83
SFMP-16-bit 53.41 81.19 82.15 78.99 81.39 72.93 75.01
2.5 AMQ 34.89 59.63 65.57 57.18 71.00 63.61 58.65
SFMP 41.13 66.58 73.49 64.35 74.05 66.46 64.34
Ours 45.90 73.40 78.38 69.02 75.57 69.14 68.57
3 AMQ 45.48 72.69 76.48 70.38 77.64 70.01 68.78
SFMP 47.40 75.80 77.43 71.89 78.13 67.80 69.74
Ours 49.06 76.60 79.88 74.75 78.18 71.35 71.64
W3G128 51.88 78.79 81.44 75.87 78.18 70.88 72.84
3.5 AMQ 49.57 77.10 80.00 76.15 79.54 73.01 72.56
SFMP 49.49 76.72 81.16 76.95 79.87 73.64 72.97
Ours 52.82 80.18 81.35 76.94 79.16 73.16 73.94
4 AMQ 50.68 78.20 81.04 77.83 79.92 73.09 73.46
SFMP 52.22 79.50 81.53 77.76 81.01 73.95 74.33
Ours 54.18 81.82 81.93 78.38 79.60 74.51 75.07
W4G128 54.01 81.65 82.42 78.84 80.03 73.40 75.06
Table 10: Detailed accuracies on six tasks for Qwen3-8B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Grayed rows denote uniform 3-bit and 4-bit baselines quantized with SignRoundV2.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 56.14 80.77 86.61 74.98 77.97 67.96 74.07
SFMP-16-bit 56.65 80.85 86.64 74.93 77.47 68.66 74.20
2.5 AMQ 35.75 56.19 75.90 55.76 69.70 58.41 58.62
SFMP 44.37 72.05 83.09 61.46 72.69 63.30 66.16
Ours 44.97 72.81 81.77 62.30 73.29 65.75 66.82
3 AMQ 47.61 73.78 84.40 66.71 73.94 63.93 68.40
SFMP 54.18 79.08 84.56 70.10 75.24 65.00 71.36
Ours 53.58 78.70 86.12 68.74 76.61 69.53 72.21
W3G128 55.55 80.30 85.72 71.50 77.69 68.67 73.24
3.5 AMQ 51.02 77.06 86.40 71.42 76.93 67.08 71.65
SFMP 55.12 78.28 85.69 72.70 76.12 68.51 72.74
Ours 54.10 80.05 86.39 71.93 77.26 68.19 72.99
4 AMQ 53.92 78.49 85.29 73.64 77.25 67.27 72.64
SFMP 55.15 79.04 85.88 74.20 77.09 68.35 73.29
Ours 56.31 79.67 86.61 73.77 77.53 68.67 73.76
W4G128 57.42 80.26 86.88 73.79 77.26 69.06 74.11
Table 11: Detailed accuracies on six tasks for Qwen3-14B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Grayed rows denote uniform 3-bit and 4-bit baselines quantized with SignRoundV2.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 60.49 82.74 89.39 78.80 80.03 73.01 77.41
SFMP-16-bit 60.41 82.79 89.33 78.92 79.98 72.84 77.38
2.5 AMQ 44.18 70.47 84.39 64.31 72.09 63.90 66.56
SFMP 50.68 75.21 86.73 69.14 76.44 68.51 71.12
Ours 51.62 78.58 87.98 70.21 77.09 70.32 72.63
3 AMQ 50.27 75.89 85.33 71.16 75.94 69.34 71.32
SFMP 57.58 80.89 87.80 75.35 78.45 71.74 75.30
Ours 57.08 80.26 87.37 74.15 77.58 71.51 74.66
W3G128 59.56 82.15 88.01 76.12 79.27 73.24 76.39
3.5 AMQ 58.31 81.56 87.56 76.04 79.12 71.98 75.76
SFMP 59.98 82.49 88.96 77.35 79.60 72.93 76.89
Ours 59.04 81.94 88.99 77.05 79.22 71.67 76.32
4 AMQ 59.42 82.05 88.76 77.68 79.65 72.13 76.62
SFMP 60.41 83.08 89.02 78.23 79.60 72.45 77.13
Ours 59.81 82.58 88.90 78.19 79.71 73.09 77.05
W4G128 61.01 82.91 89.08 78.17 80.09 72.69 77.33
Table 12: Detailed accuracies on six tasks for Qwen3-32B with mixed-bit weight-only quantization. “Avg. Bits” indicates the average bit-width. Gray rows denote uniform-bit baselines.
Avg. bits Method ARC-c ARC-e BoolQ Hella. PIQA Wino. AVG
16 Ours 60.92 83.29 86.42 82.60 81.99 73.09 78.05
SFMP-16-bit 60.92 83.25 86.42 82.56 81.88 72.93 77.99
2.5 AMQ 50.63 73.90 80.08 71.68 75.74 64.19 69.37
SFMP 55.89 78.41 82.26 76.93 79.16 67.32 73.33
Ours 55.97 79.21 88.41 74.50 78.24 73.40 74.96
3 AMQ 58.83 79.62 83.26 77.10 77.14 68.15 74.02
SFMP 59.64 81.35 86.70 80.00 79.27 70.48 76.24
Ours 58.79 80.81 89.02 78.29 79.49 73.24 76.61
W3G128 60.32 82.74 85.20 78.66 79.92 73.09 76.66
3.5 AMQ 59.71 81.15 84.78 80.02 79.14 71.26 76.01
SFMP 60.41 82.02 85.88 81.18 81.42 72.53 77.24
Ours 62.37 82.91 85.14 80.57 80.58 72.61 77.36
4 AMQ 60.87 82.31 85.42 81.58 80.95 71.76 77.15
SFMP 61.09 83.46 86.20 82.01 81.73 72.83 77.89
Ours 61.01 82.11 85.05 80.98 81.39 74.35 77.48
W4G128 61.43 82.37 85.75 82.24 81.28 73.40 77.75
Appendix C Mixed MXFP Results
C.1 Mixed MXFP4/6 Results
This subsection reports the detailed results for mixed MXFP4/6 quantization. We list the task-wise accuracies across 10 tasks for each model provided in Tables 13, 14, 15, 16, and 17.
Table 13: The detailed accuracies of 10 tasks for Llama-3.1-8B-Instruct model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
16 16-bit 51.54 81.78 84.04 59.13 73.14 67.96 33.00 80.03 37.21 73.80 64.16
4 RTN 44.37 75.55 80.86 55.19 61.96 57.06 31.80 76.28 29.25 70.80 58.31
SRV1 46.25 79.00 83.09 55.28 69.63 60.55 31.80 77.75 33.41 70.48 60.72
Ours 50.68 80.05 83.12 56.28 69.05 61.69 32.40 78.13 32.19 71.11 61.47
4.5 DL 48.38 79.29 83.88 58.31 68.48 64.51 31.60 78.78 35.01 71.74 62.00
Ours 50.43 80.13 84.86 57.83 71.36 65.75 34.60 79.43 36.60 72.45 63.34
5 DL 50.94 81.06 85.11 58.90 68.64 66.47 35.00 79.54 35.25 74.35 63.53
Ours 51.11 81.52 84.68 58.77 71.76 66.81 34.60 79.43 36.35 74.98 64.00
Table 14: The detailed accuracies of 10 tasks for Llama-3.1-70B-Instruct model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
16 16-bit 62.46 86.87 87.80 65.15 75.51 82.34 37.20 83.30 40.64 78.69 70.00
4 RTN 59.56 85.10 88.81 63.16 73.36 79.03 36.20 82.10 40.88 79.40 68.76
SRV1 58.28 85.06 88.53 63.35 75.24 79.66 37.60 82.32 40.15 79.95 69.01
Ours 59.39 85.02 89.27 63.12 74.75 79.92 37.60 82.21 40.64 81.14 69.31
4.5 DL 61.01 86.32 88.96 65.09 75.53 81.01 37.40 83.03 41.86 81.53 70.17
Ours 59.98 85.94 89.24 65.10 76.98 81.21 36.80 83.03 41.00 83.82 70.31
5 DL 61.86 86.57 88.81 65.57 76.79 81.75 37.40 83.35 42.59 82.16 70.69
Ours 61.35 86.36 88.75 65.89 76.98 82.19 38.00 83.35 40.76 82.95 70.66
Table 15: The detailed accuracies of 10 tasks for Qwen2.5-7B-Instruct model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
16 16-bit 52.56 81.61 86.36 62.01 69.44 71.76 34.60 79.65 47.98 70.72 65.67
4 RTN 48.81 75.59 82.48 56.84 57.68 65.15 34.80 75.79 41.62 67.48 60.62
SRV1 53.24 80.89 85.47 57.88 64.41 68.05 35.00 77.48 46.76 71.43 64.06
Ours 51.28 80.81 85.72 58.16 66.47 68.49 34.20 77.64 45.29 69.14 63.72
4.5 DL 50.60 80.13 86.51 59.27 69.22 69.73 33.80 76.66 45.53 70.17 64.16
Ours 53.92 81.73 86.21 59.56 69.03 70.43 34.60 77.20 45.41 69.61 64.77
5 DL 52.47 81.78 86.73 60.66 68.41 70.32 36.80 77.69 46.39 70.96 65.22
Ours 53.50 82.07 86.33 60.75 69.36 70.77 35.00 78.07 46.88 70.80 65.35
Table 16: The detailed accuracies of 10 tasks for Qwen3-8B model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
16 16-bit 55.46 83.38 86.73 57.04 64.00 72.96 31.20 76.61 36.72 68.27 63.24
4 RTN 50.17 76.14 84.98 52.00 58.08 65.90 28.60 73.01 33.66 62.83 58.54
SRV1 51.28 81.02 85.20 52.86 58.26 68.51 31.40 75.03 32.44 66.54 60.25
Ours 53.92 81.65 85.50 52.75 62.14 69.41 30.40 75.19 35.62 67.88 61.45
4.5 DL 52.05 80.30 85.96 53.96 62.57 70.99 31.00 74.32 34.27 68.03 61.35
Ours 54.95 82.66 86.67 54.82 64.37 71.52 30.80 76.12 37.33 68.19 62.74
5 DL 51.54 80.22 87.25 55.43 62.41 71.96 31.80 76.17 35.62 68.19 62.06
Ours 54.44 82.53 86.97 55.75 64.27 71.90 31.80 76.33 35.62 68.51 62.81
Table 17: The detailed accuracies of 10 tasks for Qwen3-32B model with MXFP4/6 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL”: DeltaLoss without tuning.
16 16-bit 57.94 84.47 86.39 63.91 67.13 80.74 36.00 80.96 39.05 73.40 67.00
4 RTN 56.66 81.23 84.95 61.73 65.96 76.59 33.80 77.97 39.17 70.64 64.87
SRV1 57.51 84.55 87.13 61.61 68.83 78.32 36.20 79.71 41.37 73.95 66.92
Ours 55.97 82.70 85.87 61.23 68.79 78.41 34.60 78.62 40.15 72.69 65.90
4.5 DL 57.08 83.25 86.39 63.07 66.08 79.48 33.40 80.96 42.11 70.80 66.26
Ours 56.57 84.22 88.04 63.44 68.33 79.91 36.00 80.14 40.88 73.48 67.10
5 DL 57.85 84.22 85.20 63.35 67.53 79.84 35.60 79.27 39.53 71.67 66.41
Ours 60.15 84.22 87.25 63.38 68.45 80.15 35.40 80.09 39.53 73.32 67.19
C.2 Mixed MXFP4/8 Results
This subsection reports the detailed results for mixed MXFP4/8 quantization. Table 18 reports the average accuracies and recovery rates across 10 tasks, for MXFP4/8 mixed-precision quantization. The results show that our method generally improves over RTN and achieves competitive or better accuracy than SignRoundV1 (SRV1) in the 4-bit setting. At higher average bit widths (4.5–6 bits), DeltaLoss-only (DL) and our method approach full-precision accuracy, with our method achieving the best recovery in most scenarios. Overall, these results demonstrate that our approach effectively preserves model performance under extreme low-bit mixed-precision quantization while minimizing accuracy loss. The corresponding task-wise accuracies are provided in Tables 19, 20, 21, 22, and 23.
Table 18: Average accuracies and recovery rates (%) across 10 tasks, with MXFP4 at 4 bits and mixed MXFP4/8 for the higher-bit settings. “Avg. Bits” indicates the average per-weight bit-width. “DL” denotes DeltaLoss without tuning. The “I” in the model name is short for “Instruct”.
16 16-bit 64.16 70.00 65.67 63.24 67.00
@@ C.2 Mixed MXFP4/8 Results
Table 19: The detailed accuracies of 10 tasks for Llama-3.1-8B-Instruct model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
16 16-bit 51.54 81.78 84.04 59.13 73.14 67.96 33.00 80.03 37.21 73.80 64.16
4 RTN 44.37 75.55 80.86 55.19 61.96 57.06 31.80 76.28 29.25 70.80 58.31
SRV1 46.25 79.00 83.09 55.28 69.63 60.55 31.80 77.75 33.41 70.48 60.72
Ours 47.35 79.63 82.60 56.10 69.53 62.00 33.60 77.69 32.80 72.06 61.34
Ours 49.06 80.43 83.18 57.35 71.41 64.86 33.40 78.40 34.76 70.40 62.33
5 DL 50.43 79.17 84.10 58.22 68.62 64.77 34.60 78.62 34.52 72.61 62.57
Ours 49.83 80.93 84.19 57.92 72.19 65.43 35.80 79.22 35.37 71.03 63.19
6 DL 51.45 81.19 84.59 58.94 69.73 65.98 36.40 79.33 34.76 74.03 63.64
Ours 51.54 81.40 84.89 58.57 71.96 66.85 35.40 79.38 36.84 74.35 64.12